### v1

# A topological constraint on clean Lagrangian intersections from $\mathbb{Q}$-valued augmentations

Abstract
For a knot $K$ in $\mathbb{R}^{3}$, consider its conormal bundle $L_{K}$ in $T^{*}\mathbb{R}^{3}$. We prove that if $K$ is the connected sum $K^{\prime}\#3_{1}$ of an arbitrary knot $K^{\prime}$ and the trefoil knot $3_{1}$ (or its mirror), then there is no compactly supported Hamiltonian diffeomorphism $\varphi$ on $T^{*}\mathbb{R}^{3}$ such that $\varphi(L_{K})$ and the zero section $\mathbb{R}^{3}$ intersect cleanly along the unknot in $\mathbb{R}^{3}$. For the proof, we utilize the knot DGA, whose homology was proved to be isomorphic to the knot contact homology by Ekholm, Etnyre, Ng and Sullivan, with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. We discover a certain algebraic constraint imposed on any augmentation of the knot DGA of $K=K^{\prime}\#3_{1}$ to a field $\mathbb{k}$. This constraint is non-trivial for $\mathbb{k}=\mathbb{Q}$ and plays a key role in the proof, but trivial if $\mathbb{k}$ is quadratically closed or if an augmentation factors through the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$.
1 Introduction
Background. Let us consider the cotangent bundle $T^{*}\mathbb{R}^{3}$ of $\mathbb{R}^{3}$ with a canonical symplectic structure. We identify the image of the zero section with $\mathbb{R}^{3}$. Given any knot $K$ in $\mathbb{R}^{3}$, its conormal bundle is defined by
$L_{K}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
It is a Lagrangian submanifold of $T^{*}\mathbb{R}^{3}$ and intersects $\mathbb{R}^{3}$ cleanly along the knot $K$. Let $\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ denote the set of Hamiltonian diffeomorphisms on $T^{*}\mathbb{R}^{3}$ associated to compactly supported Hamiltonians on $T^{*}\mathbb{R}^{3}$. Using these notations, we ask a question which was observed by the author in [17].
Question 1.1.
Is there $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ and $\mathbb{R}^{3}$ have a clean intersection along a knot in $\mathbb{R}^{3}$ whose isotopy class is different from $K$?
We remark that it is essential that $\varphi$ is a Hamiltonian diffeomorphism. Indeed, for any pair of knots $(K_{0},K_{1})$ in $\mathbb{R}^{3}$, there exists a $C^{\infty}$ isotopy $(\varphi^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ with compact support such that $\varphi^{1}(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. See [17, Proposition 4.1].
As a special case, the answer to Question 1.1 when $K$ is the unknot is ‘no’. This can be deduced from the results by Ganatra-Pomerleano [10, Proposition 6.29], Smith-Wemyss [21, Proposition 4.10] or Asplund-Li [1, Theorem 1.3] about clean Lagrangian intersections in the plumbings of two copies of the cotangent bundle of $S^{3}$ along the unknots. These results are concerned with Smith’s question on the ‘persistence (or rigidity) of unknottedness of Lagrangian intersections’ [10, Question 1.5]. For its relation to Question 1.1, see [17, Section 1].
In a general case, under the assumption that $\varphi(L_{K})$ for $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ and $\mathbb{R}^{3}$ intersect cleanly along a knot in $\mathbb{R}^{3}$, the author proved [17, Theorem 1.1] by using the (framed) knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$, which is a differential graded algebra (DGA) defined by Ng [14]. [17, Theorem 1.1] is stated by using the augmentation variety in $(\mathbb{C}^{*})^{2}$ derived from the set of augmentations of the knot DGA to $\mathbb{C}$. This theorem leads us to the same answer to Question 1.1 as above when $K$ is the unknot [17, Theorem 1.3] and we can also deduce some constraints on the knot type of $\varphi(L_{K})\cap\mathbb{R}^{3}$ when $K$ has certain non-trivial knot types [17, Theorem 1.4, 1.5]. However, we cannot deduce any constraint from [17, Theorem 1.1] on the knot type of $K$ if we suppose that $\varphi(L_{K})\cap\mathbb{R}^{3}$ is the unknot in $\mathbb{R}^{3}$. For the reason, see Remark 1.5 below.
Main results. Let $3_{1}$ denote the right-handed trefoil knot in $\mathbb{R}^{3}$ and $\overline{3_{1}}$ denote its mirror. The goal of this paper is to prove the next theorem which cannot be deduced from [17, Theorem 1.1].
Theorem 1.1 (Theorem 3.10).
Let $K$ be the connected sum of an arbitrary knot in $\mathbb{R}^{3}$ with $3_{1}$ or $\overline{3_{1}}$. Then, there is no $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K})$ intersects $\mathbb{R}^{3}$ cleanly along the unknot in $\mathbb{R}^{3}$.
For the proof, the geometric arguments in the present paper are similar to those in [17], which are outlined as follows: Let $K_{0}$ and $K_{1}$ be knots in $\mathbb{R}^{3}$. For $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ intersect cleanly along $K_{1}$. We use the results by Ekholm-Etnyre-Ng-Sullivan [7, 4] which show that the Chekanov-Eliashberg DGA of the unit conormal bundle of $K_{i}$ is stable tame isomorphic to the knot DGA of $K_{i}$ for $i=0,1$. The knot DGA, introduced by Ng [12, 13], is defined combinatorially from a braid representation of a knot. We then apply a DGA map associated to an exact Lagrangian cobordism between the unit conormal bundles of $K_{0}$ and $K_{1}$. This is a part of the symplectic field theory introduced by Eliashberg-Givental-Hofer [9]. Consequently, we obtain a DGA map from the knot DGA of $K_{0}$ to that of $K_{1}$. (Precisely, we need a base change on the coefficient ring of a DGA.)
The feature of the present paper is that we utilize the knot DGA, which we denote here by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ for an oriented knot $K$, with coefficients in the ring
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\cong\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})].$
Here, $S^{*}\mathbb{R}^{3}$ denotes the unit cotangent bundle of $\mathbb{R}^{3}$ and $\Lambda_{K}$ denotes the unit conormal bundle of $K$. For details, we refer to [15, 4, 16]. We note that in [17], referring to [14, 7], the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]\cong\mathbb{Z}[H_{1}(\Lambda_{K})]$ was used.
To prove Theorem 1.1, we show the following algebraic property of the knot DGA under an assumption on the isotopy class of $K$. Let $K^{\prime}\#3_{1}$ denote the connected sum of a knot $K^{\prime}$ in $\mathbb{R}^{3}$ and the trefoil knot $3_{1}$.
Lemma 1.2 (Lemma 3.9).
Let $\mathbb{k}$ be a field and $K^{\prime}$ be a knot. Then, for any augmentation $\varepsilon\colon\mathcal{A}_{*}^{\mathrm{knot}}(K^{\prime}\#3_{1})\to\mathbb{k}$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
Here, $\varepsilon(\mu)$ and $\varepsilon(U)$ are abbreviations of $\varepsilon(\mu\cdot\mathbb{1})$ and $\varepsilon(U\cdot\mathbb{1})$ for the unit $\mathbb{1}$ of the knot DGA. In fact, $D_{\varepsilon}$ is the discriminant of a quadratic equation in $\mathbb{k}$ determined by $\varepsilon$ (see the proof of Lemma 3.9). Note that if $\varepsilon$ factors through the knot DGA over $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ obtained by setting $U=1$, then $\varepsilon(U)=1$ and we have an identity
$(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)=(\varepsilon(\mu)-1+\varepsilon(\mu)^{-1})^{2},$
so the constraint on $D_{\varepsilon}$ is trivial.
For the proof of Theorem 1.1, we use the constraint from Lemma 1.2 in the case of $\mathbb{k}=\mathbb{Q}$, that is, $D_{\varepsilon}$ is the square of some rational number for any augmentation $\varepsilon$ of the knot DGA of $K^{\prime}\#3_{1}$ to $\mathbb{Q}$. (In fact, for this purpose, we may choose other fields such as $\mathbb{Q}[\sqrt{-1}]$ and $\mathbb{F}_{5}$. See Remark 3.11.) It is crucial that $\mathbb{Q}$ is not quadratically (and algebraically) closed, and we remark that the constraint on $D_{\varepsilon}$ is trivial for any $\varepsilon$ if $\mathbb{k}=\mathbb{C}$.
Remark 1.3.
In order to study the knot DGA of a knot $K$ (or the Chekanov-Eliashberg DGA of $\Lambda_{K}$), the set of augmentations to $\mathbb{k}$ has been used with $\mathbb{k}=\mathbb{C}$ or a finite field $\mathbb{F}_{p}$ for a prime number $p$. For examples of applications, see Ng [14, Section 5.2 and 5.3] and Cornwell [2] about the augmentation polynomial when $\mathbb{k}=\mathbb{C}$ and [14, Section 5.1] about the augmentation number when $\mathbb{k}=\mathbb{F}_{p}$. On the other hand, the rational numbers $D_{\varepsilon}$ from Lemma 1.2 when $\mathbb{k}=\mathbb{Q}$ are derived in a way different from those previous researches.
Lastly, we provide a theorem which is a refinement of [17, Theorem 1.1]. For an oriented knot $K$ in $\mathbb{R}^{3}$, let $V_{K}$ denote the augmentation variety defined by [16, Definition 5.1] as a subset of $(\mathbb{C}^{*})^{3}$. Let $(x,y,Z)$ be the coordinate of $\mathbb{C}^{3}$.
Theorem 1.4 (Corollary 3.5).
For oriented knots $K_{0}$ and $K_{1}$ in $\mathbb{R}^{3}$, suppose that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ intersects $\mathbb{R}^{3}$ cleanly along $K_{1}$. Then, there exist integers $b,c$ and $d$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b}Z^{c},yZ^{d},Z).$
Remark 1.5.
Identifying $\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\mid Z=1\}$ with $(\mathbb{C}^{*})^{2}$, the augmentation variety considered in [17] coincides with $V_{K}\cap\{Z=1\}\subset(\mathbb{C}^{*})^{2}$. It follows from Theorem 1.4 that there exists an injective map
$V_{K_{1}}\cap\{Z=1\}\hookrightarrow V_{K_{0}}\cap\{Z=1\}\colon(x,y)\mapsto(xy^{b},y)$
for some $b\in\mathbb{Z}$. This can be seen as a slight refinement of [17, Theorem 1.1], but it gives no constraint on $K_{0}$ if we suppose that $K_{1}$ is the unknot. The reason is that $V_{K_{1}}\cap\{Z=1\}=\{x=1\}\cup\{y=1\}$ if $K_{1}$ is the unknot (see [16, Example 3.13]) and $\{x=1\}\cup\{y=1\}$ is a subset of $V_{K_{0}}\cap\{Z=1\}$ for any knot $K_{0}$ (see [14, Proposition 5.6] and [16, Exercise 3.17]).
Notations. In this paper, for any manifold $X$, we take $\mathbb{Z}$ as the coefficient ring of the singular homology group of $X$ and denote it by $H_{*}(X)$, unless otherwise specified. Likewise, the relative homology group over $\mathbb{Z}$ is denoted by $H_{*}(X,A)$ for any submanifold $A$ of $X$.
Organization of paper. In Section 2, for a contact manifold $Y$ and its Legendrian submanifold $\Lambda$ satisfying certain conditions, we introduce the Chekanov-Eliashberg DGA of $(Y,\Lambda)$ with coefficients in $\mathbb{Z}[H_{2}(Y,\Lambda)]$. We then define a DGA map associated to an exact Lagrangian cobordism in ×$\mathbb{R}\times Y$. In Section 2.5, we will see how to apply this DGA map to a setting with a clean Lagrangian intersection as in Question 1.1. In Section 3, we combine the argument in Section 2.5 with the results of [4, 16] about the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$. Theorem 1.4 is proved in Section 3.1. Theorem 1.1 is proved by using Lemma 1.2 (Lemma 3.9) in Section 3.2. In Appendix A, by using Floer cohomology group over $\mathbb{F}_{2}[\lambda^{\pm}]$, we prove Proposition 2.7 which gives a homological constraint on $1$-dimensional clean intersections. This enhances some results in [17]. Nevertheless, the present proof is much simpler than that in [17].
Acknowledgements. The author would like to thank Kei Irie for making valuable comments. He would also like to thank Fumikazu Nagasato and Kouki Sato for fruitful discussion before this work and for giving some insight from knot theory. This work was supported by JSPS KAKENHI Grant Number JP23KJ1238.
2 Symplectic field theory for exact Lagrangian cobordisms
We fix a $2n$-dimensional simply connected manifold $P$ with a Liouville $1$-form $\lambda$ such that $P$ is the completion of a compact Liouville domain $(\bar{P},\left.\lambda\right|_{\bar{P}})$ satisfying $c_{1}(T\bar{P})=0$. Let us consider
$\begin{array}[]{ll}Y\coloneqq P\times\mathbb{R},&\alpha\coloneqq dz-\lambda,\end{array}$
where $z$ is the coordinate of $\mathbb{R}$. Then, $\alpha$ is a contact form on $Y$ and the Reeb vector field on $Y$ with respect to $\alpha$ agrees with $\partial_{z}$. Let $\pi_{P}\colon Y\to P$ denote the projection to $P$.
The arguments from Section 2.1 to Section 2.4 are completely parallel to those in [17, Section 3], except the difference of the coefficient rings. In Section 2.5, we will discuss a particular case related to clean Lagrangian intersections.
2.1 Preliminaries
We fix a complex structure $j_{0}$ on ×$\mathbb{R}\times[0,1]$ such that $(j_{0})_{(s,t)}(\partial_{s})=\partial_{t}$, where $(s,t)$ is the coordinate of ×$\mathbb{R}\times[0,1]$. Let $D\coloneqq\{z\in\mathbb{C}\mid|z|\leq 1\}$. For any $m\in\mathbb{Z}_{\geq 0}$, we fix $p_{0},p_{1},\dots,p_{m}\in\partial D$ ordered counterclockwise and define $D_{m+1}\coloneqq D\setminus\{p_{0},p_{1},\dots,p_{m}\}$. Let $\mathcal{C}_{m+1}$ denote the space of conformal structures on $D_{m+1}$ which can be extended smoothly on $D$. Given $\kappa\in\mathcal{C}_{m+1}$, we let $j_{\kappa}$ denote a complex structure on $D_{m+1}$ which associates $\kappa$. Near each boundary puncture of the Riemann surface $(D_{m+1},j_{\kappa})$, we fix a strip coordinate by a biholomorphic map
$\psi_{0}$ $\colon[0,\infty)\times[0,1]\to D_{m+1},$
$\psi_{k}$ $\colon(-\infty,0]\times[0,1]\to D_{m+1}\text{ for }k=1,\dots,m,$
with respect to $j_{0}$ and $j_{\kappa}$ such that $\lim_{s\to\infty}\psi_{0}(s,t)=p_{0}$ and $\lim_{s\to-\infty}\psi_{k}(s,t)=p_{k}$ uniformly on $t\in[0,1]$.
Let $\overline{D_{m+1}}$ be a compactification of $D_{m+1}$ obtained by gluing $[0,\infty]\times[0,1]$ via $\psi_{0}$ and $[-\infty,0]\times[0,1]$ via $\psi_{k}$ for every $k=1,\dots,m$. In a natural way, $\overline{D_{m+1}}$ has a structure of a compact oriented surface with boundary and corner.
Let $\Lambda$ be a compact Legendrian submanifold of $(Y,\alpha)$. Later, we will impose the following conditions on $\Lambda$:
  1. 1.
    $\Lambda$ is a connected and oriented manifold and admits a spin structure.
  2. 2.
    For the immersion $\left.\pi_{P}\right|_{\Lambda}\colon\Lambda\to P$, the self-intersection of $\pi_{P}(\Lambda)$ consists only of transversal double points.
  3. 3.
    For the immersed Lagrangian submanifold $\pi_{P}(\Lambda)$ of $P$, its Maslov class vanishes.
The second condition is satisfied for generic $\Lambda$. We denote the set of Reeb chords of $\Lambda$ by
${\mathcal{R}(\Lambda)\coloneqq\{a\colon[0,T]\to Y\mid T>0,\ \frac{da}{dt}(t)=(\partial_{z})_{a(t)}\text{ for every }t\in[0,T]\text{ and }a(0),a(T)\in\Lambda\}.}$
Hereafter, for any Reeb chord $(a\in[0,T]\to Y)\in\mathcal{R}(\Lambda)$, we write $T=\int a^{*}\alpha$ as $T_{a}$. For each $a\in\mathcal{R}(\Lambda)$, we fix a continuous map $w_{a}\colon\overline{D_{1}}\to Y$ such that $w_{a}(\partial D_{1})\subset\Lambda$ and
$\lim_{s\to\infty}w_{a}\circ\psi_{0}(s,t)=a(T_{a}t)$
for every $t\in[0,1]$. When the second condition above is satisfied, the Conley-Zehnder index $\mu(a)\in\mathbb{Z}$ of $a$ is defined by using the capping path $\left.w_{a}\right|_{\partial D_{1}}$ for $a$ as described in [17, Section 3.1]. We then define
$|a|\coloneqq\mu(a)-1.$
2.2 Moduli space of pseudo-holomorphic curves with boundary in an exact Lagrangian cobordism
Let $\Lambda_{+}$ and $\Lambda_{-}$ be two compact Legendrian submanifolds of $(Y,\alpha)$. Recall that a submanifold $L$ of ×$\mathbb{R}\times Y$ is called an exact Lagrangian cobordism from $\Lambda_{-}$ to $\Lambda_{+}$ if the following are satisfied:
  • There exist $r_{+},r_{-}\in\mathbb{R}$ with $r_{-}<r_{+}$ such that
    $\begin{array}[]{ll}L\cap([r_{+},\infty)\times Y)=[r_{+},\infty)\times\Lambda_{+},&L\cap((-\infty,r_{-}]\times Y)=(-\infty,r_{-}]\times\Lambda_{-},\end{array}$
    and $L\cap([r_{-},r_{+}]\times Y)$ is compact.
  • There exists a $C^{\infty}$ function $f\colon L\to\mathbb{R}$ such that $df=\left.(e^{r}\alpha)\right|_{L}$ and
    $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}\colon(-\infty,r_{-}]\times\Lambda_{-}\to\mathbb{R}$
    is a constant function. Here, $r$ is the $\mathbb{R}$-coordinate of ×$\mathbb{R}\times Y$.
In the rest of this paper, we only consider the case where $\Lambda_{-}$ is connected. Therefore, we do not need to care about the last condition that $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}$ is a constant function since
$\left.df\right|_{(-\infty,r_{-}]\times\Lambda_{-}}=\left.(e^{r}\alpha)\right|_{(-\infty,r_{-}]\times\Lambda_{-}}=0,$
which implies that $\left.f\right|_{(-\infty,r_{-}]\times\Lambda_{-}}$ is locally constant.
Associated to $L$, we use the following notations for maps between pairs of spaces:
$\begin{split}i_{+}&\colon(Y,\Lambda_{+})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{+},x),\\ i_{-}&\colon(Y,\Lambda_{-})\to(\mathbb{R}\times Y,L)\colon x\mapsto(r_{-},x),\end{split}$ (1)
and
$R_{L}\colon(\mathbb{R}\times Y,L)\to([r_{-},r_{+}]\times Y,L\cap([r_{-},r_{+}]\times Y))$
is a strong deformation retract such that $R_{L}(r,x)=(r_{+},x)$ if $r\geq r_{+}$ and $R_{L}(r,x)=(r_{-},x)$ if $r\leq r_{-}$. In addition, for every $s\in\mathbb{R}$, we define a map
$\tau_{s}\colon\mathbb{R}\times Y\to\mathbb{R}\times Y\colon(r,x)\mapsto(r+s,x).$
Let $J$ be an almost complex structure on $P$ such that $d\lambda(\cdot,J\cdot)$ is a Riemannian metric on $P$. We associate to $J$ an almost complex structure $\widetilde{J}$ on ×$\mathbb{R}\times Y$ determined by $\tau_{s}^{*}\widetilde{J}=\widetilde{J}$ for every $s\in\mathbb{R}$, $\widetilde{J}(\operatorname{Ker}\alpha)=\operatorname{Ker}\alpha$$(\pi_{P})_{*}(\widetilde{J}(v))=J((\pi_{P})_{*}(v))$ for every $v\in\operatorname{Ker}\alpha$ and $\widetilde{J}(\partial_{z})=\partial_{r}$.
For any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$, we define
^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$
to be the space of pairs $(u,\kappa)$ of $\kappa\in\mathcal{C}_{m+1}$ and a $C^{\infty}$ map $u\colon D_{m+1}\to P$ such that:
  • $du+\widetilde{J}\circ du\circ j_{\kappa}=0$.
  • $u(\partial D_{m+1})\subset L$.
  • There exist $s_{0},s_{1},\dots,s_{m}\in\mathbb{R}$ such that
    $\lim_{s\to\infty}\tau_{-s}\circ u\circ\psi_{0}$ $=(s_{0},a(T_{a}(1-t))),$
    $\lim_{s\to-\infty}\tau_{s}\circ u\circ\psi_{k}$ $=(s_{k},b_{k}(T_{b_{k}}(1-t)))\text{ for }k=1,\dots,m,$
    $C^{\infty}$-uniformly on $[0,1]$.
  • Let $\bar{u}\colon\overline{D_{m+1}}\to L$ denote the continuous extension of the map $R_{L}\circ u\colon D_{m+1}\to L$. Then, the 2-chain
    $\bar{u}+(i_{+}\circ w_{a})-((i_{-}\circ w_{b_{1}})+\dots+(i_{-}\circ w_{b_{m}}))$
    represents the homology class $C$. (More precisely, each term is regarded as a singular 2-chain after taking suitable triangulations of $\overline{D_{m+1}}$ and $\overline{D_{1}}$.)
When $m\leq 1$, $\mathcal{C}_{m+1}$ consists of a single element and the group $\mathcal{A}_{m+1}$ of conformal automorphisms on $D_{m+1}$ acts on ^$\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})$ by $(u,\kappa)\cdot\varphi\coloneqq(u\circ\varphi,\kappa)$ for every $\varphi\in\mathcal{A}_{m+1}$. As in [17, (10) in Section 3.2], let us define a moduli space to be
$\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})\coloneqq\begin{cases}\hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})/\mathcal{A}_{m+1}&\text{ if }m=0,1,\\ \hat{\mathcal{M}}_{L,C}(a;b_{1},\dots,b_{m})&\text{ if }m\geq 2.\end{cases}$
In the case where $\Lambda_{+}=\Lambda_{-}=\Lambda$ and $L=\mathbb{R}\times\Lambda$, for any $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$, an $\mathbb{R}$-action on ×$\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m})$ is defined by $s\cdot(u,\kappa)\coloneqq(\tau_{s}\circ u,\kappa)$ for every $s\in\mathbb{R}$. Let us denote the quotient space by
¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\coloneqq\mathbb{R}\backslash\mathcal{M}_{\mathbb{R}\times\Lambda,(i_{+})_{*}(A)}(a;b_{1},\dots,b_{m}).$
2.3 Chekanov-Eliashberg DGA over $\mathbb{Z}[H_{2}(Y,\Lambda)]$
We consider a compact Legendrian submanifold $\Lambda$ of $(Y,\alpha)$ satisfying the three conditions in Section 2.1. We also assume that $\Lambda$ is admissible in the sense of the definition in [6, Section 2.3]. Let us take a spin structure $\mathfrak{s}$ on $\Lambda$.
We take an almost complex structure $J$ on $P$ as in [6, Proposition 2.3] such that the following hold for every $a,b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda)$ and $A\in H_{2}(Y,\Lambda)$: The moduli space ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is cut out transversely and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)-1$. Moreover, when $|b_{1}|+\dots+|b_{m}|=|a|-1$, $\coprod_{A\in H_{2}(Y,\Lambda)}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. (For the relation between $J$-holomorphic disks in $P$ defined as in [6, Section 2.3] and $\widetilde{J}$-holomorphic maps in ×$\mathbb{R}\times Y$ defined as in Section 2.2, see [3, Theorem 2.1] and [17, Appendix A.4].)
In addition, ¯$\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$ has an orientation depending on $\mathfrak{s}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda$. See [17, Definition A.4] for details which follow the conventions in [5] related to orientations.
Let $\mathcal{A}_{*}(\Lambda)$ be the graded $\mathbb{Z}[H_{2}(Y,\Lambda)]$-algebra freely generated by $\mathcal{R}(\Lambda)$ whose graded is given by $|\cdot|\colon\mathcal{R}(\Lambda)\to\mathbb{Z}$. A derivative over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of degree $-1$
$\partial_{J,\mathfrak{s}}\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*-1}(\Lambda)$
is defined for every $a\in\mathcal{R}(\Lambda)$ by
$\partial_{J,\mathfrak{s}}(a)\coloneqq\sum_{|b_{1}|+\dots+|b_{m}|=|a|-1}\sum_{A\in H_{2}(Y,\Lambda)}(-1)^{(n-1)(|a|+1)}\left(\#_{\operatorname{sign}}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})\right)e^{A}\cdot b_{1}\cdots b_{m}$
and extended by the Leibniz rule. Here, #$\#_{\operatorname{sign}}$ is the number of the points counted with the signs induced by the orientation.
We refer to [5, Theorem 4.1, Theorem 4.12] which takes $\mathbb{Z}[H_{1}(\Lambda)]$ as the coefficient ring of the Chekanov-Eliashberg DGA. We modify the arguments in their proofs to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda)]$ as the coefficient ring. Then, we have the following theorem.
Theorem 2.1.
$\partial_{J,\mathfrak{s}}$ is a differential, that is, $\partial_{J,\mathfrak{s}}\circ\partial_{J,\mathfrak{s}}=0$. Moreover, the stable tame isomorphism class over $\mathbb{Z}[H_{2}(Y,\Lambda)]$ of the DGA $(\mathcal{A}_{*}(\Lambda),\partial_{J,\mathfrak{s}})$ is independent of the choice of $J$ and invariant under Legendrian isotopies of $\Lambda$ preserving spin structures.
Let us discuss the effect by changing the spin structure on $\Lambda$. For any two spin structures $\mathfrak{s}_{0}$ and $\mathfrak{s}_{1}$ on $\Lambda$, a cohomology class $d(\mathfrak{s}_{0},\mathfrak{s}_{1})\in H^{1}(\Lambda;\mathbb{Z}/2)$ is defined as in [5, Subsection 4.4.1]. Let $\partial\colon H_{2}(Y,\Lambda)\to H_{1}(\Lambda)$ denote the boundary map. For any $\beta\in H_{1}(\Lambda)$, if we change the choice of a spin structure from $\mathfrak{s}$ to $\mathfrak{s}^{\prime}$, the orientation on the moduli space
$\coprod_{|b_{1}|+\dots+|b_{m}|=|a|-1}\coprod_{A\in H_{2}(Y,\Lambda)\colon\partial A=\beta}\bar{\mathcal{M}}_{\Lambda,A}(a;b_{1},\dots,b_{m})$
is changed by the sign $(-1)^{\langle d(\mathfrak{s},\mathfrak{s}^{\prime}),\beta\rangle}$. See the proof of [5, Theorem 4.29]. We define a ring isomorphism
$\psi(\mathfrak{s},\mathfrak{s}^{\prime})\colon\mathbb{Z}[H_{2}(Y,\Lambda)]\to\mathbb{Z}[H_{2}(Y,\Lambda)]\colon e^{A}\mapsto(-1)^{\langle d(\mathfrak{s},\mathfrak{s}),\partial A\rangle}e^{A}.$ (2)
Then, the next proposition easily follows as in the proof of [5, Theorem 4.30].
Proposition 2.2.
We define a graded ring isomorphism $I(\mathfrak{s},\mathfrak{s}^{\prime})\colon\mathcal{A}_{*}(\Lambda)\to\mathcal{A}_{*}(\Lambda)$ by
$I(\mathfrak{s},\mathfrak{s}^{\prime})(k\cdot a)=\psi(\mathfrak{s},\mathfrak{s}^{\prime})(k)\cdot a$
for every $k\in\mathbb{Z}[H_{2}(Y,\Lambda)]$ and $a\in\mathcal{R}(\Lambda)$. Then $\partial_{J,\mathfrak{s}^{\prime}}\circ I(\mathfrak{s},\mathfrak{s}^{\prime})=I(\mathfrak{s},\mathfrak{s}^{\prime})\circ\partial_{J,\mathfrak{s}}$.
2.4 DGA map associated to an exact Lagrangian cobordism
We consider an exact Lagrangian cobordism $L$ from $\Lambda_{-}$ to $\Lambda_{+}$ as in Section 2.2 such that the Maslov class of $L$ vanishes and both $\Lambda_{+}$ and $\Lambda_{-}$ are connected. We assume that $L$ is oriented and has a spin structure $\mathfrak{s}_{L}$. Then, on a boundary $\Lambda_{+}$ (resp. $\Lambda_{-}$) of $L$, an orientation and a spin structure $\mathfrak{s}_{+}$ (resp. $\mathfrak{s}_{-}$) are induced. See [17, Remark 3.6] for their precise descriptions.
Associated to $(\Lambda_{+},\mathfrak{s}_{+})$ and $(\Lambda_{-},\mathfrak{s}_{-})$, the Chekanov-Eliashberg DGAs
$\begin{array}[]{cc}(\mathcal{A}_{*}(\Lambda_{+}),\partial_{J,\mathfrak{s}_{+}}),&(\mathcal{A}_{*}(\Lambda_{-}),\partial_{J,\mathfrak{s}_{-}})\end{array}$
are defined as in Section 2.3. The maps $i_{+}$ and $i_{-}$ of (1) induce ring homomorphisms
$\begin{array}[]{ll}(i_{+})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{+})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)],&(i_{-})_{*}\colon\mathbb{Z}[H_{2}(Y,\Lambda_{-})]\to\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)].\end{array}$
Via $(i_{-})_{*}$, we obtain a DGA over $\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]$
$(\mathcal{A}_{*}(\Lambda_{-})\otimes_{\mathbb{Z}[H_{2}(Y,\Lambda_{-})]}\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)],\partial_{J,\mathfrak{s}_{-}}\otimes\operatorname{id}_{\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]})$
and we regard it as a DGA over $\mathbb{Z}[H_{2}(Y,\Lambda_{+})]$ via $(i_{+})_{*}$.
By [8, Lemma 3.8], the moduli space $\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is cut out transversely for generic $J$ and it is a manifold of dimension $|a|-(|b_{1}|+\dots+|b_{m}|)$ for any $a\in\mathcal{R}(\Lambda_{+})$, $b_{1},\dots,b_{m}\in\mathcal{R}(\Lambda_{-})$ and $C\in H_{2}(\mathbb{R}\times Y,L)$. By the Gromov compactness, if $|b_{1}|+\dots+|b_{m}|=|a|$, $\coprod_{C\in H_{2}(\mathbb{R}\times Y,L)}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})$ is a compact 00-dimensional manifold. In addition, it has an orientation depending on $\mathfrak{s}_{L}$ and a choice of an orientation of the determinant line of the capping operator for each Reeb chord of $\Lambda_{+}$ and $\Lambda_{-}$. See [17, Section A.3].
We refer to the proof of [17, Theorem 3.7] which takes $\mathbb{Z}[H_{1}(\Lambda_{\pm})]$ as the coefficient rings. We modify the arguments in [17, Appendix A.6] to the present case where we take $\mathbb{Z}[H_{2}(Y,\Lambda_{\pm})]$ as the coefficient rings. (As referred in [17, Appendix A], see also [11, Theorem 2.5] about the DGA map over $\mathbb{Z}$, but note that some conventions in [11] related to orientations are different from those in [5, 17].) Then, we obtain the following theorem.
Theorem 2.3.
We define a graded $\mathbb{Z}[H_{2}(Y,\Lambda_{+})]$-algebra map
$\Phi_{L}\colon\mathcal{A}_{*}(\Lambda_{+})\to\mathcal{A}_{*}(\Lambda_{-})\otimes_{\mathbb{Z}[H_{2}(Y,\Lambda_{-})]}\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]$
such that for every $a\in\mathcal{R}(\Lambda_{+})$,
$\Phi_{L}(a)=\sum_{|b_{1}|+\dots+|b_{m}|=|a|}\sum_{C\in H_{2}(\mathbb{R}\times Y,L)}(-1)^{(n-1)(|a|+1)+m}\left(\#_{\operatorname{sign}}\mathcal{M}_{L,C}(a;b_{1},\dots,b_{m})\right)b_{1}\cdots b_{m}\otimes e^{C}.$
Then, $\Phi_{L}$ is a DGA map, that is, $\Phi_{L}\circ\partial_{J,\mathfrak{s}_{+}}=(\partial_{J,\mathfrak{s}_{-}}\otimes\operatorname{id}_{\mathbb{Z}[H_{2}(\mathbb{R}\times Y,L)]})\circ\Phi_{L}$.
2.5 Application to clean Lagrangian intersections
2.5.1 Chekanov-Eliashberg DGA of the unit conormal bundle of a knot in $\mathbb{R}^{3}$
Let $(q_{1},q_{2},q_{3})$ be the coordinate on $\mathbb{R}^{3}$. The manifold $\mathbb{R}^{3}$ is equipped with a standard Riemannian metric $\langle\cdot,\cdot\rangle\coloneqq\sum_{i=1}^{3}dq_{i}\otimes dq_{i}$ and an orientation such that $dq_{1}\wedge dq_{2}\wedge dq_{3}$ is a positive volume form. A coordinate $((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))$ on $T^{*}\mathbb{R}^{3}$ is defined via the diffeomorphism
$\mathbb{R}^{3}\times\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto\sum_{i=1}^{3}p_{i}(dq_{i})_{(q_{1},q_{2},q_{3})}.$
For any $r>0$, we denote $D^{*}_{r}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|\leq r\}$. The canonical Liouville form on $T^{*}\mathbb{R}^{3}$ is given by $\lambda_{\mathbb{R}^{3}}=\sum_{i=1}^{3}p_{i}dq_{i}$.
We consider the unit cotangent bundle of $\mathbb{R}^{3}$
$S^{*}\mathbb{R}^{3}\coloneqq\{(q,p)\in T^{*}\mathbb{R}^{3}\mid|p|=1\}.$
It is endowed with a contact form $\alpha_{\mathbb{R}^{3}}\coloneqq\left.\lambda_{\mathbb{R}^{3}}\right|_{S^{*}\mathbb{R}^{3}}$. Note that there exists a diffeomorphism
$S^{*}\mathbb{R}^{3}\to T^{*}S^{2}\times\mathbb{R}\colon(q,p)\mapsto((p,q-\langle q,p\rangle p),\langle q,p\rangle)$
for the unit sphere $S^{2}\subset\mathbb{R}^{3}$ such that $\alpha_{\mathbb{R}^{3}}$ on $S^{*}\mathbb{R}^{3}$ agrees with the pullback of the $1$-form $dz-\lambda_{S^{2}}$ on $T^{*}S^{2}\times\mathbb{R}$, where $\lambda_{S^{2}}$ is the canonical Liouville form on $T^{*}S^{2}$. Therefore, by taking $(T^{*}S^{2},\lambda_{S^{2}})$ as $(P,\lambda)$, we can define the Chekanov-Eliashberg DGA for Legendrian submanifolds of $S^{*}\mathbb{R}^{3}$ as in Section 2.3.
We focus on a special kind of Legendrian submanifolds. For any knot $K$ in $\mathbb{R}^{3}$, recall the conormal bundle $L_{K}$ defined in Section 1. Then, the unit conormal bundle of $K$ is defined by
$\Lambda_{K}\coloneqq L_{K}\cap S^{*}\mathbb{R}^{3}=\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ \langle p,v\rangle=0\text{ for every }v\in T_{q}K\}.$
It is a Legendrian submanifold of $S^{*}\mathbb{R}^{3}$.
Let us prepare several notations. We identify a disk bundle of $L_{K}$ with a tubular neighborhood of $K$ in $\mathbb{R}^{3}$ via the map $L_{K}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\mathbb{R}^{3}\colon(q,p)\mapsto q+p$ for sufficiently small $\varepsilon>0$. Then, an orientation on $L_{K}$ is determined from the orientation on $\mathbb{R}^{3}$. We also fix a spin structure $\mathfrak{s}_{L_{K}}$ given by a trivial $\mathrm{Spin}(3)$-bundle on $L_{K}$. As the positive end of $L_{K}$, an orientation and a spin structure $\mathfrak{s}_{K}$ on $\Lambda_{K}$ are determined as in [17, Remark 3.6].
We fix an orientation on $K$. Let $l\colon K\to\Lambda_{K}$ be a section which gives the zero framing of $K$ and $m\colon S^{1}\to\Lambda_{K}$ be the meridian of $K$. The conventions are as follows: $S^{1}$ is oriented as the boundary of the unit disk $D$, and $(\nu,dl(v_{l}),dm(v_{m}))$ is a positive basis of $T_{q}\mathbb{R}^{3}$ at $q\in l(K)\cap m(S^{1})$, where $\nu\in T_{q}(L_{K}\cap D^{*}_{1}\mathbb{R}^{3})$ points outward and $v_{l}$ and $v_{m}$ are positive vectors tangent to $K$ and $S^{1}$ respectively. Then, the homology classes $[l]$ and $[m]$ give a basis of $H_{1}(\Lambda_{K})$.
We choose a non-vanishing vector field $V$ on $K$. Then, $\Lambda_{K}$ is bounded by a compact orientable manifold
$\{(q,p)\in S^{*}\mathbb{R}^{3}\mid q\in K,\ p(V_{q})\geq 0\}.$
Therefore, the homology class $[\Lambda_{K}]\in H_{2}(S^{*}\mathbb{R}^{3})$ vanishes. For the pair $(S^{*}\mathbb{R}^{3},\Lambda_{K})$, we have a short exact sequence
[…]
We choose a basis $\{\widehat{l},\widehat{m},\widehat{s}\}$ of $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ such that $\partial(\widehat{l})=[l]$, $\partial(\widehat{m})=[m]$ and $\widehat{s}$ is represented by the fiber $S^{*}_{q}\mathbb{R}^{3}$ of any $q\in\mathbb{R}^{3}$. For this choice, we refer to the remark after the proof of Theorem 1.2 in [4, Section 3.7], which means that we use a knot isotopic to $K$ and transversal to the contact structure $\operatorname{Ker}(dq_{3}-q_{1}dq_{2}+q_{2}dq_{1})$ on $\mathbb{R}^{3}$.
Remark 2.4.
From the above short exact sequence, the lifts of $[l]$ and $[m]$ to $H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})$ can only be well-defined modulo $\mathbb{Z}\cdot\widehat{s}$. For the proof of the main results in Section 3, it is not important how to specify the lifts $\widehat{l}$ and $\widehat{m}$ (see Theorem 3.2).
We denote $\lambda\coloneqq e^{\widehat{l}}$, $\mu\coloneqq e^{\widehat{m}}$ and $U\coloneqq e^{\widehat{s}}$, then
$\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K})]=\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}].$
From the arguments in Section 2.3, the Chekanov-Eliashberg DGA $(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}_{K}})$ for $(\Lambda_{K},\mathfrak{s}_{K})$ is defined with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$.
2.5.2 Exact Lagrangian cobordism arising from clean intersection
Given a $C^{\infty}$ Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support, we define $(\varphi^{t}_{H})_{t\in[0,1]}$ to be the flow of a time-dependent vector field $(X_{H}^{t})_{t\in[0,1]}$ on $T^{*}\mathbb{R}^{3}$ characterized by $d\lambda_{\mathbb{R}^{3}}(\cdot,X^{t}_{H})=d(H(\cdot,t))$ for every $t\in[0,1]$. We denote
$\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})\coloneqq\{\varphi=\varphi^{1}_{H}\mid H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}\text{ is a compactly supported }C^{\infty}\text{ Hamiltonian}\}.$
The image of the zero section of $T^{*}\mathbb{R}^{3}$ is identified with $\mathbb{R}^{3}$.
Let us recall the definition of a clean intersection of submanifolds.
Definition 2.5.
Let $L$ and $L^{\prime}$ be submanifolds of $T^{*}\mathbb{R}^{3}$. We say that $L$ and $L^{\prime}$ have a clean intersection along $L\cap L^{\prime}$ if the following hold:
  • $L\cap L^{\prime}$ is a submanifold of both $L$ and $L^{\prime}$.
  • For every $x\in L\cap L^{\prime}$, $T_{x}(L\cap L^{\prime})=T_{x}L\cap T_{x}L^{\prime}$.
Let us make a setup to deal with Question 1.1. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$ and tak $\varphi\in\mathrm{Ham}_{c}(T^{*}\mathbb{R}^{3})$. Suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. From [17, Lemma 2.3], we may assume that there exists $\varepsilon>0$ such that
$\varphi(L_{K_{0}})\cap D_{\varepsilon}^{*}\mathbb{R}^{3}=L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$
by moving $\varphi(L_{K_{0}})$ by a Hamiltonian isotopy supported in a neighborhood of $\varphi(L_{K_{0}})\cap\mathbb{R}^{3}=K_{1}$. We consider a diffeomorphism
$F\colon\mathbb{R}\times S^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\setminus\mathbb{R}^{3}\colon(r,(q,p))\mapsto(q,e^{r}p),$
for which $F^{*}(\lambda_{\mathbb{R}^{3}})=e^{r}\alpha_{\mathbb{R}^{3}}$. Furthermore,
$L_{\varphi}\coloneqq F^{-1}(\varphi(L_{K_{0}})\setminus K_{1})$
is an exact Lagrangian cobordism from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$. The inclusion maps (1) are written as
$i_{+}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{+},(q,p)),$
$i_{-}\colon(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})\to(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})\colon(q,p)\mapsto(r_{-},(q,p)),$
for some $r_{+},r_{-}\in\mathbb{R}$ with $r_{+}>r_{-}$.
The Lagrangian filling $\varphi(L_{K_{0}})$ of $\Lambda_{K_{0}}$ in $T^{*}\mathbb{R}^{3}$ is obtained by gluing the Lagrangian cobordism $L_{\varphi}$ in ×$\mathbb{R}\times S^{*}\mathbb{R}^{3}$ and the Lagrangian filling $L_{K_{1}}$ in $T^{*}\mathbb{R}^{3}$ along $\Lambda_{K_{1}}$. Therefore, as a part of the Mayer-Vietoris exact sequence, we have a short exact sequence
[…]
Note that $H_{2}(T^{*}\mathbb{R}^{3},L_{K_{1}})\cong H_{1}(L_{K_{1}})=\mathbb{Z}\cdot[K_{1}]$ and $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))\cong H_{1}(L_{K_{0}})=\mathbb{Z}\cdot[K_{0}]$. The left map is a surjection since $(i_{+})_{*}(\widehat{l})\in H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is mapped to a generator of $H_{2}(T^{*}\mathbb{R}^{3},\varphi(L_{K_{0}}))$. It follows that $H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})$ is a free abelian group of rank $3$ with a basis $\{\widehat{l}_{0},\widehat{m}_{0},\widehat{s}_{0}\}$ defined by
$\widehat{l}_{0}$ $\coloneqq(i_{+})_{*}(\widehat{l})\text{ for }\widehat{l}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}}),$
$\widehat{m}_{0}$ $\coloneqq(i_{-})_{*}(\widehat{m})\text{ for }\widehat{m}\in H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}}),$
$\widehat{s}_{0}$ [$\coloneqq[\{0\}\times S^{*}_{q}\mathbb{R}^{3}]\text{ for any }q\in\mathbb{R}^{3}.$
We abbreviate
$\begin{array}[]{ccc}R_{+}\coloneqq\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{0}})],&R_{-}\coloneqq\mathbb{Z}[H_{2}(S^{*}\mathbb{R}^{3},\Lambda_{K_{1}})],&R_{0}\coloneqq\mathbb{Z}[H_{2}(\mathbb{R}\times S^{*}\mathbb{R}^{3},L_{\varphi})].\end{array}$
The maps $i_{+}$ and $i_{-}$ induce ring homomorphisms $(i_{+})_{*}\colon R_{+}\to R_{0}$ and $(i_{-})_{*}\colon R_{-}\to R_{0}$. Let us denote
$\begin{array}[]{ccc}\lambda_{0}\coloneqq e^{\widehat{l}_{0}}=(i_{+})_{*}(\lambda),&\mu_{0}\coloneqq e^{\widehat{m}_{0}}=(i_{-})_{*}(\mu),&U_{0}\coloneqq e^{\widehat{s}_{0}}=(i_{+})_{*}(U)=(i_{-})_{*}(U).\end{array}$
Then $R_{0}=\mathbb{Z}[(\lambda_{0})^{\pm},(\mu_{0})^{\pm},(U_{0})^{\pm}]$.
Lemma 2.6.
$(i_{+})_{*}\colon R_{+}\to R_{0}$ and $(i_{-})_{*}\colon R_{-}\to R_{0}$ are ring isomorphisms. Moreover, There exist $a\in\{\pm 1\}$ and $b,c,d\in\mathbb{Z}$ such that the isomorphism $(i_{-})_{*}^{-1}\circ(i_{+})_{*}$ is given by
$(i_{-})_{*}^{-1}\circ(i_{+})_{*}\colon R_{+}\to R_{-}\colon\begin{cases}\lambda\mapsto\lambda^{a}\mu^{b}U^{c},\\ \mu\mapsto\mu^{a}U^{d},\\ U\mapsto U.\end{cases}$ (3)
To prove this lemma, we use the following result.
Proposition 2.7.
The open embedding $\varphi^{-1}\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ preserves the orientations of $L_{K_{1}}$ and $L_{K_{0}}$. Moreover, $\varphi_{*}([K_{0}])=a\cdot[K_{0}]$ in $H_{1}(\varphi(L_{K_{0}}))$ for $a\in\{1,-1\}$.
The proof is left to Appendix A. Note that the first assertion about orientations was shown in [17, Lemma 4.4] and a version of the second assertion with coefficients in $\mathbb{F}_{2}$ was shown by [17, Proposition 2.8, Lemma 4.5] using Floer cohomology group over $\mathbb{F}_{2}$ with the triangle product. It should be possible to prove Proposition 2.7 in a way parallel to [17, Section 2.3] by lifting the coefficient ring from $\mathbb{F}_{2}$ to $\mathbb{Z}$. However, we will give a simpler proof in Appendix A by using Floer cohomology with coefficients in $\mathbb{F}_{2}[H_{1}(L_{K_{0}})]\cong\mathbb{F}_{2}[\lambda^{\pm}]$.
Proof of Lemma 2.6.
It is clear that $(i_{\pm})_{*}\colon R_{\pm}\to R_{0}$ maps $U$ to $U_{0}$. The circle $\varphi(K_{0})$ is oriented via $\varphi\colon K_{0}\to\varphi(K_{0})$. From Proposition 2.7, we take $a\in\{\pm 1\}$ such that $[K_{1}]=a[\varphi(K_{0})]$. Consider the map induced by $i_{\pm}$ on the first homology groups
$\begin{array}[]{ll}(i_{+})_{*}\colon H_{1}(\Lambda_{K_{0}})\to H_{1}(L_{\varphi}),&(i_{-})_{*}\colon H_{1}(\Lambda_{K_{1}})\to H_{1}(L_{\varphi}).\end{array}$
A basis of $H_{1}(L_{\varphi})$ is given by $\{l_{0},m_{0}\}$ for $l_{0}\coloneqq(i_{+})_{*}([l])$ and $m_{0}\coloneqq(i_{-})_{*}([m])$. It suffices to show that for some $b\in\mathbb{Z}$,
$\begin{array}[]{cc}(i_{-})_{*}([l])=a\cdot l_{0}+b\cdot m_{0},&(i_{+})_{*}([m])=a\cdot m_{0}.\end{array}$ (5)
By the map $H_{1}(L_{\varphi})\to H_{1}(\varphi(L_{K_{0}}))$ induced by $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})$, $l_{0}$ is mapped to $[\varphi(K_{0})]$ and $m_{0}$ is mapped to 00. Since $(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([l])=[K_{1}]=a[\varphi(K_{0})]$ in $H_{1}(\varphi(L_{K_{0}}))$, the coefficient of $l_{0}$ in $(i_{-})_{*}([l])$ is $a$. This shows the first equation of (5) for some $b\in\mathbb{Z}$.
Next, take an embedded disk $D\subset\varphi(L_{K_{0}})$ which bounds the circle $(F\circ i_{+}\circ m)(S^{1})$. $D$ is oriented so that
$[\partial D]=(\left.F\right|_{L_{\varphi}})_{*}((i_{+})_{*}([m]))\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1})$
for $\left.F\right|_{L_{\varphi}}\colon L_{\varphi}\to\varphi(L_{K_{0}})\setminus K_{1}$. We may assume that $D$ intersects $K_{1}$ transversely. Then, for sufficiently small $\varepsilon>0$, $D\cap(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})=\coprod_{i\in I}D^{\prime}_{i}$, where $D^{\prime}_{i}\subset D$ is a small disk and $I$ is a finite set. Then, using $a\in\{\pm 1\}$ determined by $[K_{1}]=a[\varphi(K_{0})]$, we have
$\sum_{i\in I}[\partial D_{i}]=a\cdot(\left.F\right|_{L_{\varphi}}\circ i_{-})_{*}([m])=a\cdot(\left.F\right|_{L_{\varphi}})_{*}(m_{0})\text{ in }H_{1}(\varphi(L_{K_{0}})\setminus K_{1}).$
For the first equality, note that the inclusion map $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to\varphi(L_{K_{0}})$ preserves orientations by Proposition 2.7 when $\varphi(L_{K_{0}})$ is oriented via the map $\varphi\colon L_{K_{0}}\to\varphi(L_{K_{0}})$. Let $\Sigma$ be the closure of $F^{-1}(D\setminus\coprod_{i\in I}D^{\prime}_{i})$ in $L_{\varphi}$. Then,
$\begin{array}[]{cc}0=[\partial\Sigma]=(\left.F\right|_{L_{\varphi}})_{*}^{-1}([\partial D])-\sum_{i\in I}(\left.F\right|_{L_{\varphi}})^{-1}_{*}([\partial D_{i}])=(i_{+})_{*}([m])-a\cdot m_{0}&\text{in }H_{1}(L_{\varphi})\end{array}$
This shows the second equation of (5). ∎
The exact Lagrangian cobordism $L_{\varphi}$ from $\Lambda_{K_{1}}$ to $\Lambda_{K_{0}}$ induces a DGA map as below.
Proposition 2.8.
There exists a DGA map over $R_{+}$
$\Phi_{L_{\varphi}}\colon(\mathcal{A}_{*}(\Lambda_{K_{0}}),\partial_{J,\mathfrak{s}_{K_{0}}})\to(\mathcal{A}_{*}(\Lambda_{K_{1}}),\partial_{J,\mathfrak{s}_{K_{1}}}),$
where the $R_{-}$-algebra $\mathcal{A}_{*}(\Lambda_{K_{1}})$ is changed to an $R_{+}$-algebra via the isomorphism (3).
Proof.
$L_{K_{0}}$ is equipped with the spin structure $\mathfrak{s}_{L_{K_{0}}}$ given by a trivial $\mathrm{Spin}(3)$-bundle. It is restricted on $L_{K_{0}}\setminus\varphi^{-1}(K_{1})$, then via the diffeomorphism $F^{-1}\circ\varphi\colon L_{K_{0}}\setminus\varphi^{-1}(K_{1})\to L_{\varphi}$, we obtain a spin structure $s_{L_{\varphi}}$ on the exact Lagrangian cobordism $L_{\varphi}$. This induces $\mathfrak{s}_{K_{0}}$ on the positive end $\Lambda_{K_{0}}$ and $\mathfrak{s}_{K_{1}}$ on the negative end $\Lambda_{K_{0}}$. (Here, we need to care about the underlying orientations of spin manifolds. See the arguments in [17, Subsection 4.2.3].) Then, the existence of a DGA map $\Phi_{L_{\varphi}}$ follows from Theorem 2.3 and the description of the base change from $R_{-}$ to $R_{+}$ follows from Lemma 2.6. ∎
Lastly, let us introduce a spin structure $\mathfrak{s}^{0}_{K}$ on $\Lambda_{K}$ as in [7, Section 6.2]. When $T\Lambda_{K}$ is trivialized as an oriented vector bundle, then $\mathfrak{s}^{0}_{K}$ is given by the trivial $\mathrm{Spin}(2)$-bundle on $\Lambda_{K}$. It determines the orientations of moduli spaces which are used to define the differential of the DGAs in [4, Section 2.3.4] and [7, Section 3.6]. The ring isomorphism $\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})$ from (2) is given by
$\psi(\mathfrak{s}^{0}_{K},\mathfrak{s}_{K})\colon\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\to\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]\colon\lambda\mapsto\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$ (6)
3 Proof of main results
In this section, we abbreviate the Laurent polynomial ring $\mathbb{Z}[\lambda^{\pm},\mu^{\pm},U^{\pm}]$ by $R$.
3.1 General case
We will combine Proposition 2.8 with the results in [4, 15, 16] about the knot DGA. Some conventions about the knot DGA in these papers are slightly different from each other, as summarized in [16, Appendix A]. In this section, we mainly refer to [16] for the definition of the knot DGA.
Let $K$ be an oriented knot in $\mathbb{R}^{3}$ given by the closure of a braid $B$. Then, a DGA over $R$, called the knot DGA of $K$, is constructed from $B$ as in [16, Definition 3.11]. Its stable tame isomorphism class as a DGA over $R$ depends only on the isotopy class of $K$ as an oriented knot. See [15, Theorem 1.4] and [16, Theorem 3.1]. Let us denote this DGA by $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ and its homology $\operatorname{Ker}\partial_{K}/\operatorname{Im}\partial_{K}$ by $HC_{*}(K)$. We will review a description of the 00-th degree part $HC_{0}(K)$ in Section 3.2.
Theorem 3.1 (Theorem 1.2 of [4], Theorem 3.14 and Appendix A of [16] ).
$(\mathcal{A}_{*}(\Lambda_{K}),\partial_{J,\mathfrak{s}^{0}_{K}})$ is stable tame isomorphic over $R$ to $(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})$ whose $R$-algebra structure is changed via a ring isomorphism
$\psi^{\prime}\colon R\to R\colon\lambda\mapsto-\lambda,\ \mu\mapsto-\mu,\ U\mapsto U.$
The base change by $\psi^{\prime}$ is due to the difference of conventions between [4] and [16]. See [16, Appendix A]. It follows from this theorem that there are unital graded ring homomorphisms
$\begin{array}[]{cc}I_{K}\colon\mathcal{A}_{*}(\Lambda_{K})\to\mathcal{A}^{\mathrm{knot}}_{*}(K),&\bar{I}_{K}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K)\to\mathcal{A}_{*}(\Lambda_{K}),\end{array}$ (8)
intertwining $\partial_{K}$ and $\partial_{J,\mathfrak{s}^{0}_{K}}$ such that
$\begin{array}[]{cc}I_{K}(r\cdot x)=\psi^{\prime}(r)\cdot I_{K}(x),&\bar{I}_{K}(r\cdot y)=(\psi^{\prime})^{-1}(r)\cdot\bar{I}_{K}(y),\end{array}$
for every $r\in R$, $x\in\mathcal{A}_{*}(\Lambda_{K})$ and $y\in\mathcal{A}^{\mathrm{knot}}_{*}(K)$.
We continue to consider the setup in Subsection 2.5.2. Let $K_{0}$ and $K_{1}$ be oriented knots in $\mathbb{R}^{3}$. For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, suppose that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along $K_{1}$. Then we have the following theorem analogous to [17, Proposition 4.6].
Theorem 3.2.
There exist integers $a,b,c$ and $d$ with $a\in\{\pm 1\}$ and a unital graded ring homomorphism
$\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})$
such that $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ and
$\Psi_{\varphi}(r\cdot x)=\psi(r)\cdot\Psi_{\varphi}(x)$ (9)
for every $x\in\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})$ and $r\in R$, where $\psi$ a ring isomorphism defined by
$\psi\colon R\to R\colon\lambda\mapsto\lambda^{a}\mu^{b}U^{c},\ \mu\mapsto\mu^{a}U^{d},\ U\mapsto U.$
Proof.
A unital graded ring homomorphism $\Psi_{\varphi}$ satisfying $\Psi_{\varphi}\circ\partial_{K_{0}}=\partial_{K_{1}}\circ\Psi_{\varphi}$ is defined as the composition of the following morphisms of differential graded rings:
[…]
Here, $\bar{I}_{K_{0}}$ and $I_{K_{1}}$ are the morphisms (8) from Theorem 3.1. $I(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})$ for $i\in\{0,1\}$ are the isomorphisms in Proposition 2.2. The vertical morphism $\Phi_{L_{\varphi}}$ is given by Proposition 2.8. From the relation between each morphism and the $R$-algebra structure on each DGA, the relation (9) holds if we define $\psi$ by
[…]
where $\psi^{\prime}$ is defined in Theorem 3.1 and $\psi(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})$ for $i\in\{0,1\}$ are given by (6). We compute that for $i\in\{0,1\}$
$\psi(\mathfrak{s}^{0}_{K_{i}},\mathfrak{s}_{K_{i}})\circ(\psi^{\prime})^{-1}\colon\lambda\mapsto-\lambda,\ \mu\mapsto\mu,\ U\mapsto U,$
and $\psi$ is described as in the theorem for the integers $a,b,c$ and $d$ which determine (3). ∎
We recall the notion of an augmentation of a DGA. See, for instance, [16, Definition 2.6].
Definition 3.3.
Let $\mathbb{k}$ be a field and $(\mathcal{A}_{*},\partial)$ be a DGA over $R$. A ring homomorphism $\varepsilon\colon\mathcal{A}_{*}\to\mathbb{k}$ is called an augmentation of $(\mathcal{A}_{*},\partial)$ to $\mathbb{k}$ if $\varepsilon(\mathbb{1})=1$, $\varepsilon(x)=0$ for any $x\in\mathcal{A}_{p}$ with $p\neq 0$, and $\varepsilon\circ\partial\colon\mathcal{A}_{1}\to\mathbb{k}$ is the zero map. Here, $\mathbb{1}$ is the unit of $\mathcal{A}_{*}$.
A knot invariant is defined from the set of augmentations of the knot DGA to $\mathbb{C}$.
Definition 3.4 (Definition 5.1 of [16]).
For an oriented knot $K$ in $\mathbb{R}^{3}$, its augmentation variety is a subset of $(\mathbb{C}^{*})^{3}$ defined by
$V_{K}\coloneqq\left\{(x,y,Z)\in(\mathbb{C}^{*})^{3}\ \middle|\begin{array}[]{l}\text{There exists an augmentation }\varepsilon\text{ of }(\mathcal{A}^{\mathrm{knot}}_{*}(K),\partial_{K})\text{ to }\mathbb{C}\\ \text{such that }\varepsilon(\lambda)=x,\ \varepsilon(\mu)=y\text{ and }\varepsilon(U)=Z\end{array}\right\}.$
Here, $\varepsilon(r\cdot\mathbb{1})$ for any $r\in R$ is abbreviated by $\varepsilon(r)$.
From Theorem 3.2, we obtain a refinement of [17, Theorem 1.1].
Corollary 3.5.
There exist integers $b^{\prime},c^{\prime}$ and $d^{\prime}$ for which we have an injective map
$V_{K_{1}}\hookrightarrow V_{K_{0}}\colon(x,y,Z)\mapsto(xy^{b^{\prime}}Z^{c^{\prime}},yZ^{d^{\prime}},Z).$
Proof.
By Theorem 3.2, for any augmentation $\varepsilon$ of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{1}),\partial_{K_{1}})$ to $\mathbb{C}$ with $(\varepsilon(\lambda),\varepsilon(\mu),\varepsilon(U))=(x,y,Z)$, the composite map $\varepsilon\circ\Psi_{\varphi}\colon\mathcal{A}_{*}^{\mathrm{knot}}(K_{0})\to\mathbb{C}$ is an augmentation of $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ to $\mathbb{C}$ such that
$(\varepsilon\circ\Psi_{\varphi}(\lambda),\varepsilon\circ\Psi_{\varphi}(\mu),\varepsilon\circ\Psi_{\varphi}(U))=(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
From the definition of $V_{K_{i}}$ for $i\in\{0,1\}$, this implies that we have a map
$V_{K_{1}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{a}y^{b}Z^{c},y^{a}Z^{d},Z).$
We use the following fact by [15, Proposition 4.3]: Up to chain isomorphisms of DGA over $R$, $(\mathcal{A}_{*}^{\mathrm{knot}}(K_{0}),\partial_{K_{0}})$ is unchanged by the substitution $\lambda\mapsto\lambda^{-1}U$, $\mu\mapsto\mu^{-1}U$ and $U\mapsto U$. Hence, we have a bijection
$V_{K_{0}}\to V_{K_{0}}\colon(x,y,Z)\mapsto(x^{-1}Z,y^{-1}Z,Z).$
If $a=1$, we take $b^{\prime}\coloneqq b$, $c^{\prime}\coloneqq c$ and $d^{\prime}\coloneqq d$. If $a=-1$, we take $b^{\prime}\coloneqq-b$, $c^{\prime}\coloneqq-c+1$ and $d^{\prime}\coloneqq-d+1$. Then, the assertion of the corollary holds. ∎
3.2 Proof of Theorem 1.1
First, let us review how to compute $HC_{0}(K)$ for an oriented knot $K$. We refer to [16, Section 3]. Note that since $\mathcal{A}^{\mathrm{knot}}_{p}(K)=0$ for $p<0$, $HC_{0}(K)=\mathcal{A}^{\mathrm{knot}}_{0}(K)/\partial_{K}(\mathcal{A}^{\mathrm{knot}}_{1}(K))$.
For any $n\in\mathbb{Z}_{\geq 1}$, let $B_{n}$ denote the group of braids on $n$ strands, which is generated by $\{\sigma_{1},\dots,\sigma_{n-1}\}$ with relations $\sigma_{i}\sigma_{i+1}\sigma_{i}=\sigma_{i+1}\sigma_{i}\sigma_{i+1}$ for $21\leq i\leq n-$ 2 and $\sigma_{i}\sigma_{j}=\sigma_{j}\sigma_{i}$ for $|i-j|\geq 2$. Let $\mathcal{A}_{n}$ be the $R$-algebra freely generated by the set $\{a_{i,j}\mid i,j\in\{1,\dots,n\},\ i\neq j\}$. Then, a group homomorphism $B_{n}\to\mathrm{Aut}(\mathcal{A}_{n})\colon B\mapsto\phi_{B}$ is defined as in [16, Definition 3.3]. It is determined by $\phi_{\sigma_{k}}\colon\mathcal{A}_{n}\to\mathcal{A}_{n}$ for each generator $\sigma_{k}$, and $\phi_{\sigma_{k}}$ is given by
$\phi_{\sigma_{k}}\colon\begin{cases}a_{i,j}\mapsto a_{i,j}&\text{if }i,j\neq k,k+1,\\ a_{k+1,i}\mapsto a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k+1}\mapsto a_{i,k}&\text{ if }i\neq k,k+1,\\ a_{k,k+1}\mapsto-a_{k+1,k},&\\ a_{k+1,k}\mapsto-a_{k,k+1},&\\ a_{k,i}\mapsto a_{k+1,i}-a_{k+1,k}a_{k,i}&\text{ if }i\neq k,k+1,\\ a_{i,k}\mapsto a_{i,k+1}-a_{i,k}a_{k,k+1}&\text{ if }i\neq k,k+1.\end{cases}$ (10)
See also [7, Section 2.1]. We remark that $\phi_{B\cdot B^{\prime}}=\phi_{B}\circ\phi_{B^{\prime}}$ for any $B,B^{\prime}\in B_{n}$.
The automorphism $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n})$ for $B\in B_{n}$ is extended to an automorphism on $\mathcal{A}_{n+1}$ as follows: We consider the group homomorphism $B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$ defined as above and an injective map $B_{n}\to B_{n+1}$ defined by adding the extra $(n+1)$-th strand to any braid on $n$ strands (see [16, Remark 3.5]). Then, for any $B\in B_{n}$, $\phi_{B}\in\mathrm{Aut}(\mathcal{A}_{n+1})$ is defined via the composite map $B_{n}\hookrightarrow B_{n+1}\to\mathrm{Aut}(\mathcal{A}_{n+1})$. In addition, let us label the $(n+1)$-st strand by * and denote $a_{i,n+1}$ by $a_{i,*}$ and $a_{n+1,i}$ by $a_{*,i}$ for every $i\in\{1,\dots,n\}$.
We define $n\times n$ matrices $\mathbb{A}$ and $\widehat{\mathbb{A}}$ by
$\begin{array}[]{ll}\mathbb{A}_{i,j}\coloneqq\begin{cases}a_{i,j}&\text{ if }i<j,\\ 1-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j,\end{cases}&\widehat{\mathbb{A}}_{i,j}\coloneqq\begin{cases}Ua_{i,j}&\text{ if }i<j,\\ U-\mu&\text{ if }i=j,\\ -\mu a_{i,j}&\text{ if }i>j.\end{cases}\end{array}$
Fix $B\in B_{n}$ which has the writhe $w$ (the sum of the exponents in the braid word). Let $\bm{\Lambda}$ be an $n\times n$ diagonal matrix with $\bm{\Lambda}_{1,1}=\lambda\mu^{w}U^{-(w-n+1)/2}$ and $\bm{\Lambda}_{i,i}=1$ for every $i\in\{2,\dots,n\}$. Let $\bm{\Phi}^{L}_{B}$ be an $n\times n$ matrix whose entries $(\bm{\Phi}^{L}_{B})_{i,j}\in\mathcal{A}_{n}$ for $i,j\in\{1,\dots,n\}$ are determined by
$\phi_{B}(a_{i,*})=\sum_{j=1}^{n}(\bm{\Phi}^{L}_{B})_{i,j}a_{j,*}.$
Let $K$ be the oriented knot in $\mathbb{R}^{3}$ given by the closure of the braid $B$. By [16, Theorem 4.1, Remark 4.2], $HC_{0}(K)$ is the quotient of $\mathcal{A}_{n}$ by the two-sided ideal generated by all entries of the two matrices
$\begin{array}[]{cc}\mathbb{A}-\bm{\Lambda}\cdot\phi_{B}(\mathbb{A})\cdot\bm{\Lambda}^{-1},&\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B}\cdot\mathbb{A},\end{array}$ (12)
where $(\phi_{B}(\mathbb{A}))_{i,j}\coloneqq\phi_{B}(\mathbb{A}_{i,j})$ for $i,j\in\{1,\dots,n\}$. The isomorphism class of $HC_{0}(K)$ as a unital $R$-algebra depends only on the isotopy class of $K$. See [16, Theorem 3.15] and [15, Theorem 1.4 (2)].
We consider a commutative $R$-algebra $HC_{0}^{\mathrm{ab}}(K)$ defined as the quotient of $HC_{0}(K)$ by the relations $\xi\eta=\eta\xi$ for all $\xi,\eta\in HC_{0}(K)$. Using the $R$-algebra $R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]$ of polynomials and a natural map $\pi_{n}\colon\mathcal{A}_{n}\to R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]\colon a_{i,j}\mapsto a_{i,j}$, we have
$HC_{0}^{\mathrm{ab}}(K)=R[a_{i,j}\mid 1\leq i,j\leq n,\ i\neq j]/\mathcal{I}_{B},$
where $\mathcal{I}_{B}$ is the ideal generated by the images under $\pi_{n}$ of all entries of the two matrices (12).
[…]
Figure. 1: The braid $B^{\#}\in B_{n+2}$. The box with ‘$B$’ in the center represents the braid $B$ on $n$ strands.
Associated to $B$, consider a braid $B^{\#}$ on $(n+2)$ strands defined by
$B^{\#}\coloneqq\sigma_{n}\cdot(\sigma_{n+1})^{3}\cdot B\in B_{n+2}.$
See Figure 1. The closure of $B^{\#}$ gives a knot which is the connected sum $K\#3_{1}$ of $K$ and the right-handed trefoil knot $3_{1}$.
In general, $HC^{\mathrm{ab}}_{0}(K\#3_{1})$ is a complicated $R$-algebra. However, it has the following property.
Lemma 3.6.
We define an $R$-coefficient polynomial $F\in R[X]$ by
$F(X)\coloneqq X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1).$
Then, there exists a unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1}).$
Proof.
We compute $\phi_{B^{\#}}(a_{n+2,*})$ by (10). Since $\phi_{\sigma_{k}}(a_{n+2,*})=a_{n+2,*}$ for every $k\in\{1,\dots,n-1\}$, $\phi_{B}(a_{n+2,*})=a_{n+2,*}$. Furthermore,
${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{B}}$${a_{n+2,*}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+2,*}-a_{n+2,n+1}a_{n+1,*}}$$\scriptstyle{\phi_{\sigma_{n+1}}}$${a_{n+1,*}+a_{n+1,n+2}(a_{n+2,*}-a_{n+2,n+1}a_{n+1,*})}$$\scriptstyle{\phi_{\sigma_{n}}}$${a_{n,*}+a_{n,n+2}(a_{n+2,*}-a_{n+2,n}a_{n,*}).}$
Let us write $P=a_{n,n+2}$ and $Q=a_{n+2,n}$. Then, the bottom $(n+2)$-th row of $\bm{\Phi}^{L}_{B^{\#}}$ is
$\begin{pmatrix}0&\cdots&0&1-PQ&0&P\end{pmatrix}.$
It follows that the $(n+2,n)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n}-((1-PQ)\mathbb{A}_{n,n}+P\mathbb{A}_{n+2,n})$ $=-\mu Q-((1-PQ)(1-\mu)+P(-\mu Q))$
$=PQ-\mu Q-(1-\mu),$
and the $(n+2,n+2)$-th entry of $\widehat{\mathbb{A}}-\bm{\Lambda}\cdot\bm{\Phi}^{L}_{B^{\#}}\cdot\mathbb{A}$ is
$\widehat{\mathbb{A}}_{n+2,n+2}-((1-PQ)\mathbb{A}_{n,n+2}+P\mathbb{A}_{n+2,n+2})$ $=(U-\mu)-((1-PQ)P+P(1-\mu))$
$=PQP-(2-\mu)P+(U-\mu).$
Hence $P^{2}Q^{2}-(2-\mu)PQ+(U-\mu)Q\in\mathcal{I}_{B^{\#}}$ and by substituting $PQ$ with $\mu Q+(1-\mu)$, we can find an element of $\mathcal{I}_{B^{\#}}$
$(\mu Q+(1-\mu))^{2}-(2-\mu)(\mu Q+(1-\mu))+(U-\mu)Q,$
which is equal to
$\mu^{2}Q^{2}+(2\mu(1-\mu)-(2-\mu)\mu+(U-\mu))Q+(1-\mu)^{2}-(2-\mu)(1-\mu)$
=\ = $\mu^{2}Q^{2}+(-\mu^{2}-\mu+U)Q+(\mu-1).$
Therefore, if we define $F(X)=X^{2}+(\mu+1-\mu^{-1}U)X+(\mu-1)$, we have a well-defined unital $R$-algebra map
$J\colon R[X]/(F)\to HC^{\mathrm{ab}}_{0}(K\#3_{1})$
which sends $X$ to $-\mu Q=-\mu a_{n+2,n}$. ∎
Let us check this property in the simplest case.
Example 3.7.
From a braid representation $\sigma_{1}^{3}\in B_{2}$ of $3_{1}$, it is not difficult to compute $HC_{0}(3_{1})$ as outlined in [16, Exercise 3.10]. $HC_{0}(3_{1})$ is isomorphic to a commutative $R$-algebra $R[Y]/(G_{1},G_{2})$, where
$\begin{array}[]{cc}G_{1}(Y)\coloneqq UY^{2}-\mu UY+\lambda\mu^{3}(1-\mu),&G_{2}(Y)\coloneqq UY^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-U).\end{array}$
See [16, Exercise 5.5]. We have identities
$YG_{1}+\mu G_{2}$ $=UY^{3}+\lambda\mu^{3}(2-\mu)Y+\lambda\mu^{3}(\mu-U),$
$UY(YG_{1}+\mu G_{2})$ $=(UY^{2})^{2}+\lambda\mu^{3}(2-\mu)(UY^{2})+\lambda\mu^{3}(\mu-U)UY,$
$UY^{2}$ $=G_{1}+(\mu UY-\lambda\mu^{3}(1-\mu)),$
so the ideal $(G_{1},G_{2})$ has an element
$(\mu UY-\lambda\mu^{3}(1-\mu))^{2}+\lambda\mu^{3}(2-\mu)(\mu UY-\lambda\mu^{3}(1-\mu))+\lambda\mu^{3}(\mu-U)UY$
=\ = $(\mu UY)^{2}+(\lambda\mu^{4}+\lambda\mu^{3}-\lambda\mu^{2}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1).$
This means that a unital $R$-algebra map
$R[X]/(F)\to HC^{\mathrm{ab}}_{0}(3_{1})\cong R[Y]/(G_{1},G_{2})\colon X\mapsto\lambda^{-1}\mu^{-2}UY=(\lambda\mu^{3})^{-1}(\mu UY)$
is well-defined for the polynomial $F$ in Lemma 3.6, since
$(\lambda\mu^{3})^{2}F(\lambda^{-1}\mu^{-2}UY)=(\mu UY)^{2}+(\lambda\mu^{3})(\mu+1-\mu^{-1}U)(\mu UY)+(\lambda\mu^{3})^{2}(\mu-1)\in(G_{1},G_{2}).$
Remark 3.8.
If we reduce the coefficient ring $R$ to $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$ by setting $U=1$, Lemma 3.6 implies the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to\left.HC^{\mathrm{ab}}_{0}(K\#3_{1})\right|_{U=1}$
induced by $J$. (Here, the notation $\left.HC_{0}(K\#3_{1})\right|_{U=1}$ for an $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra is the same one as in [16, Section 4].) However, this assertion is trivial. Indeed,
$\left.F\right|_{U=1}(X)=X^{2}+(\mu+1-\mu^{-1})X+(\mu-1)=(X+\mu)(X+1-\mu^{-1})$
and thus, for any unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra $A$, there exists a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][X]/(\left.F\right|_{U=1})\to A\colon X\mapsto-\mu$.
Lemma 3.6 implies the following constraint on augmentations.
Lemma 3.9.
Let $\mathbb{k}$ be a field. For any augmentation $\varepsilon$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ to $\mathbb{k}$,
$D_{\varepsilon}\coloneqq(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))^{2}-4(\varepsilon(\mu)-1)$
satisfies $D_{\varepsilon}\in\{k^{2}\mid k\in\mathbb{k}\}$.
Proof.
Since $HC_{0}(K\#3_{1})=\mathcal{A}^{\mathrm{knot}}_{0}(K\#3_{1})/\partial_{K\#3_{1}}(\mathcal{A}^{\mathrm{knot}}_{1}(K\#3_{1}))$, any augmentation $\varepsilon$ to $\mathbb{k}$ of $(\mathcal{A}^{\mathrm{knot}}_{*}(K\#3_{1}),\partial_{K\#3_{1}})$ induces a ring homomorphism
$\bar{\varepsilon}\colon HC^{\mathrm{ab}}_{0}(K\#3_{1})\to\mathbb{k}$
such that $\bar{\varepsilon}(r)=\varepsilon(r)$ for every $r\in R$. Composing with $J$ in Lemma 3.6, we obtain a map
$\bar{\varepsilon}\circ J\colon R[X]/(F)\to\mathbb{k}.$
Let us denote $\zeta=\bar{\varepsilon}\circ J(X)$. Since $\bar{\varepsilon}\circ J(F)=0$ and $J$ is a unital $R$-algebra map, we obtain an equation in $\mathbb{k}$
$\zeta^{2}+(\varepsilon(\mu)+1-\varepsilon(\mu)^{-1}\varepsilon(U))\zeta+(\varepsilon(\mu)-1)=0.$
Then, $D_{\varepsilon}$ in the lemma is the discriminant of this quadratic equation for $\zeta$. In order that this equation has a solution in $\mathbb{k}$, it is necessary that $D_{\varepsilon}$ is the square of some element of $\mathbb{k}$. ∎
Combining Theorem 3.2, Lemma 3.9 and a computation of the knot DGA of the unknot, we obtain a constraint on the isotopy class of clean Lagrangian intersections.
Theorem 3.10.
Let $K_{0}$ be the connected sum of an arbitrary knot $K$ in $\mathbb{R}^{3}$ with $3_{1}$ or its mirror $\overline{3_{1}}$. Then, there is no $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection along the unknot in $\mathbb{R}^{3}$.
Proof.
Assume that there exists $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ such that $\varphi(L_{K_{0}})$ and $\mathbb{R}^{3}$ have a clean intersection $K_{1}=\varphi(L_{K_{0}})\cap\mathbb{R}^{3}$ and $K_{1}$ is the unknot in $\mathbb{R}^{3}$. In the case of $K_{0}=K\#3_{1}$, we can deduce a contradiction as follows: From the trivial braid on a single strand, we can compute as in [16, Example 3.13] that $\mathcal{A}^{\mathrm{knot}}_{0}(K_{1})=R$ and $\partial_{K_{1}}(\mathcal{A}_{1}^{\mathrm{knot}}(K_{1}))$ is the ideal generated by
$U-\lambda-\mu+\lambda\mu.$
Let us choose an augmentation $\varepsilon_{1}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{1})\to\mathbb{Q}$ determined by
$\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1.\end{array}$
Composing with the map $\Psi_{\varphi}$ in Theorem 3.2, we obtain $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}\colon\mathcal{A}^{\mathrm{knot}}_{*}(K_{0})\to\mathbb{Q}$. It is an augmentation of the knot DGA of $K_{0}=K\#3_{1}$ to $\mathbb{Q}$ such that
$\begin{array}[]{ccc}\varepsilon(\lambda)=3^{a}2^{b}(-1)^{c},&\varepsilon(\mu)=2^{a}(-1)^{d},&\varepsilon(U)=-1.\end{array}$
We apply Lemma 3.9 for $\mathbb{k}=\mathbb{Q}$. Then,
$D_{\varepsilon}=(\varepsilon(\mu)+1+\varepsilon(\mu)^{-1})^{2}-4(\varepsilon(\mu)-1)$
must be the square of some rational number. However, we can calculate as follows:
$D_{\varepsilon}=\begin{cases}(\frac{7}{2})^{2}-4\cdot 1=\frac{33}{4}&\text{ if}\varepsilon(\mu)=2,\\ (-\frac{3}{2})^{2}-4\cdot(-3)=\frac{57}{4}&\text{ if }\varepsilon(\mu)=-2,\\ (\frac{7}{2})^{2}-4\cdot(-\frac{1}{2})=\frac{57}{4}&\text{ if }\varepsilon(\mu)=\frac{1}{2},\\ (-\frac{3}{2})^{2}-4\cdot(-\frac{3}{2})=\frac{33}{4}&\text{ if }\varepsilon(\mu)=-\frac{1}{2},\end{cases}$
and this is a contradiction.
In the case of $K_{0}=K\#\overline{3_{1}}$, we can deduce a contradiction as follows: Take a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ such that $\varphi=\varphi^{1}_{H}$. Consider an involution
$\iota\colon T^{*}\mathbb{R}^{3}\to T^{*}\mathbb{R}^{3}\colon(q,p)=((q_{1},q_{2},q_{3}),(p_{1},p_{2},p_{3}))\mapsto((q_{1},q_{2},-q_{3}),(p_{1},p_{2},-p_{3}))$
which preserves $\lambda_{\mathbb{R}^{3}}$. Then, $\iota\circ\varphi^{t}_{H}\circ\iota=\varphi^{t}_{\overline{H}}$ for $\overline{H}\coloneqq H\circ(\iota\times\operatorname{id}_{[0,1]})$. Moreover, for the knots
$\overline{K_{i}}\coloneqq\{(q_{1},q_{2},-q_{3})\in\mathbb{R}^{3}\mid(q_{1},q_{2},q_{3})\in K_{i}\}\text{ for }i=0,1,$
$\varphi^{1}_{\overline{H}}(L_{\overline{K_{0}}})$ and $\mathbb{R}^{3}=\iota(\mathbb{R}^{3})$ intersect cleanly along $\overline{K_{1}}$. Since $\overline{K_{1}}$ is the unknot and $\overline{K_{0}}$ is the connected sum of the mirror of $K$ and $3_{1}$, this contradicts to the result in the previous case. ∎
Remark 3.11.
For the proof of the theorem, we may choose other field $\mathbb{k}$ and other augmentation $\varepsilon_{1}$ of the knot DGA of $K\#3_{1}$ to $\mathbb{k}$ such that $\varepsilon(U)\neq 1$. Let us see two examples.
  1. 1.
    Choose $\mathbb{k}=\mathbb{Q}[\sqrt{-1}]$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=2-\sqrt{-1},&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=\sqrt{-1},\end{array}$
    in $\mathbb{Q}[\sqrt{-1}]$. For $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$, let us write $D_{\varepsilon}=D_{y}$ if $\varepsilon(\mu)=y$. Then,
    $D_{y}=(y+1-\sqrt{-1}y^{-1})^{2}-4(y-1)=y^{2}-2y-2\sqrt{-1}y^{-1}-y^{-2}+(5-2\sqrt{-1}),$
    and $y$ varies in the set {±$\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$. Noting that $D_{\sqrt{-1}y^{-1}}=D_{y}$, we can compute that $D_{y}$ takes its value in the set
    ${S\coloneqq\left\{\frac{19}{4}-3\sqrt{-1},\frac{51}{4}-\sqrt{-1},\frac{1}{4}-6\sqrt{-1},\frac{9}{4}+2\sqrt{-1}\right\}.}$
    In general, suppose that $D=a+\sqrt{-1}b$ for $a,b\in\mathbb{Q}$ satisfies $D\in\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$. Then, $a^{2}+b^{2}=D\cdot\bar{D}\in\{r^{2}\mid r\in\mathbb{Q}\}$, where $\bar{D}$ is the complex conjugate of $D$. However, for any $D=a+\sqrt{-1}b\in S$, we can check that $a^{2}+b^{2}\notin\{r^{2}\mid r\in\mathbb{Q}\}$. Therefore, $D_{y}\notin\{k^{2}\mid k\in\mathbb{Q}[\sqrt{-1}]\}$ for any $y\in\{\pm 2,\pm 2\sqrt{-1},\pm\frac{1}{2},\pm\frac{\sqrt{-1}}{2}\}$.
  2. 2.
    Choose $\mathbb{k}=\mathbb{F}_{5}\cong\mathbb{Z}/5$ and $\varepsilon_{1}$ determined by
    $\begin{array}[]{ccc}\varepsilon_{1}(\lambda)=3,&\varepsilon_{1}(\mu)=2,&\varepsilon_{1}(U)=-1,\end{array}$
    in $\mathbb{F}_{5}$, similar to the one in the proof of Theorem 3.10. In general, if $D\in\mathbb{Z}/5$ satisfies $D\equiv k^{2}\mod 5$ for some $k\in\mathbb{Z}$, then $D\equiv 0,1$ or $4\mod 5$. However, for $\varepsilon\coloneqq\varepsilon_{1}\circ\Psi_{\varphi}$,
    $D_{\varepsilon}\equiv 33\cdot 4\equiv 2\text{ or }D_{\varepsilon}\equiv 57\cdot 4\equiv 3\mod 5.$
    Therefore, $D_{\varepsilon}\notin\{k^{2}\mid k\in\mathbb{F}_{5}\}$.
Remark 3.12.
To the author’s knowledge, it seems difficult to show Theorem 3.10, even when $K_{0}=3_{1}$, by only using the knot DGA with coefficients in $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$. Indeed, for the trefoil knot $3_{1}$ and the unknot $O$,
$\left.HC_{0}(3_{1})\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}][Y]/(Y^{2}-\mu Y+\lambda\mu^{3}(1-\mu),\ Y^{2}+\lambda\mu^{2}Y+\lambda\mu^{2}(\mu-1)),$
$\left.HC_{0}(O)\right|_{U=1}$ $\cong\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]/((1-\lambda)(1-\mu)),$
and we can find a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map
$\left.HC_{0}(3_{1})\right|_{U=1}\to\left.HC_{0}(O)\right|_{U=1}\colon Y\mapsto\mu-\mu^{2}.$
Therefore, we cannot deduce a contradiction from the existence of a unital $\mathbb{Z}[\lambda^{\pm},\mu^{\pm}]$-algebra map which is induced on the 00-the degree part of the homology from the DGA map as in Theorem 3.2.
A Proof of Proposition 2.7
Let $K_{0}$ be an oriented knot in $\mathbb{R}^{3}$ and consider its conormal bundle $L_{K_{0}}$. The Lagrangian submanifolds $\mathbb{R}^{3}$ and $L_{K_{0}}$ of $T^{*}\mathbb{R}^{3}$ are oriented as in Subsection 2.5.1.
For any $\psi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we introduce a $\mathbb{Z}/2$-graded Floer cochain complex
$(CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H),d_{J})$
with coefficients in the Laurent polynomial ring $\mathbb{F}_{2}[\lambda^{\pm}]$. It is a slight enhancement of the Floer cochain complex over $\mathbb{F}_{2}$ constructed as in [20, Section 8].
We choose a Hamiltonian $H\colon T^{*}\mathbb{R}^{3}\times[0,1]\to\mathbb{R}$ with compact support such that the Lagrangian submanifold $P_{H}\coloneqq(\varphi^{1}_{H}\circ\psi)(\mathbb{R}^{3})$ intersects $L_{K_{0}}$ transversely. $P_{H}$ is oriented so that the diffeomorphism $\varphi^{1}_{H}\circ\psi\colon\mathbb{R}^{3}\to P_{H}$ preserves orientations. From the orientations of $P_{H}$ and $L_{K_{0}}$, a $\mathbb{Z}/2$-grading $|x|\in\mathbb{Z}/2$ is defined for every $x\in P_{H}\cap L_{K_{0}}$. See [19, Example 2.7, Section 2.d]. Note that the grading is shifted by $1$ if we use the opposite orientation of $P_{H}$. We also choose an almost complex structure $J$ such that $d\lambda_{\mathbb{R}^{3}}(\cdot,J\cdot)$ is a Riemannian metric on $T^{*}\mathbb{R}^{3}$ and $J(\partial_{p_{i}})=\partial_{q_{i}}$ for $i\in\{1,2,3\}$ outside a compact subset of $T^{*}\mathbb{R}^{3}$.
[…]
Figure. 2: The shaded region is the image of $u$. $c_{x}$ and $c_{y}$ are paths in $L_{K_{0}}$ and $u(\mathbb{R}\times\{1\})$ is contained in $L_{K_{0}}$. $u(\mathbb{R}\times\{0\})$ is contained in $\psi(\mathbb{R}^{3})$.
Fix a base point $z_{0}\in L_{K_{0}}$. For any $x\in P_{H}\cap L_{K_{0}}$, we choose a path $c_{x}\colon[0,1]\to L_{K_{0}}$ such that $c_{x}(0)=z_{0}$ and $c_{x}(1)=x$. Let $c_{x}^{-1}$ denote the path $[0,1]\to L_{K_{0}}\colon t\mapsto c_{x}(1-t)$. For every $x,y\in P_{H}\cap L_{K_{0}}$ and a homology class $A\in H_{1}(L_{K_{0}})$, we define $\mathcal{M}_{(H,J),A}(x,y)$ to be the space of maps $u\colon\mathbb{R}\times[0,1]\to T^{*}\mathbb{R}^{3}$ satisfying:
  • su($\partial_{s}u(s,t)+J_{u(s,t)}(\partial_{t}u(s,t)-(X^{t}_{H})_{u(s,t)})=0$ for every $(s,t)\in\mathbb{R}\times[0,1]$.
  • $u(\mathbb{R}\times\{0\})\subset\psi(\mathbb{R}^{3})$ and $u(\mathbb{R}\times\{1\})\subset L_{K_{0}}$.
  • $\lim_{s\to-\infty}u(s,t)=\varphi^{t}_{H}((\varphi^{1}_{H})^{-1}(x))$ and $\lim_{s\to\infty}u(s,t)=\varphi_{H}^{t}((\varphi^{1}_{H})^{-1}(y))$ $C^{\infty}$-uniformly on $t\in[0,1]$.
  • $A\in H_{1}(L_{K_{0}})$ is represented by a loop based at $z_{0}$ obtained by concatenating the three paths in $L_{K_{0}}$: $c_{x}$, the path $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto u(s,1)$ from $x$ to $y$, and $c_{y}^{-1}$.
Figure 2 depicts $u(\mathbb{R}\times[0,1])\subset T^{*}\mathbb{R}^{3}$ and the three paths in $L_{K_{0}}$. The group $\mathbb{R}$ acts on $\mathcal{M}_{(H,J),A}(x,y)$ by $(u\cdot\sigma)(s,t)=u(s+\sigma,t)$ for every $\sigma\in\mathbb{R}$.
To define the $\mathbb{Z}/2$-graded Floer cochain complex, we refer to [19, Section 2.f]. We take the pair $(H,J)$ to be regular so that the moduli space $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is cut out transversely. The dimension of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is equal to $|x|-|y|-1$ modulo $2$. Moreover, by the Gromov compactness, the union of 00-dimensional components of $\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ is a finite set. We define $n_{(H,J),A}(x,y)\in\mathbb{F}_{2}$ to be the modulo $2$ count of the number of the 00-dimensional components of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$.
We define $CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)$ to be the $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module freely generated by the set $P_{H}\cap L_{K_{0}}$ whose grading is given by $|\cdot|\colon P_{H}\cap L_{K_{0}}\to\mathbb{Z}/2$. The $\mathbb{F}_{2}[\lambda^{\pm}]$-linear map $d_{J}$ of degree $1$ is defined by
$d_{J}(y)\coloneqq\sum_{m=-\infty}^{\infty}\sum_{|x|\equiv|y|+1\mod 2}\left(n_{(H,J),m[K_{0}]}(x,y)\right)\lambda^{m}\cdot x$
for every $y\in P_{H}\cap L_{K_{0}}$. By a standard argument using the Gromov compactification of $\mathcal{M}_{(H,J),A}(x,y)/\mathbb{R}$ for each $A\in H_{1}(L_{K_{0}})$ when its dimension is $1$, we can prove that $d_{J}\circ d_{J}=0$.
From the cochain complex, we obtain a cohomology group
$HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H,J))\coloneqq\operatorname{Ker}d_{J}/\operatorname{Im}d_{J}.$
Its isomorphism class as a $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-module is independent of the choice of the paths $\{c_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$. To see this claim, we choose another path $c^{\prime}_{x}$ instead of $c_{x}$ for every $x\in P_{H}\cap L_{K_{0}}$ and define $d^{\prime}_{J}$ by using $\{c^{\prime}_{x}\mid x\in P_{H}\cap L_{K_{0}}\}$ in the same way as $d_{J}$. Take $m_{x}\in\mathbb{Z}$ such that $m_{x}[K_{0}]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating $c^{\prime}_{x}$ and $c_{x}^{-1}$. Then we can define an $\mathbb{F}_{2}$-linear isomorphism
$\begin{array}[]{cc}\Psi\colon CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)\to CF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};H)\colon x\mapsto\lambda^{m_{x}}x&\text{for }x\in P_{H}\cap L_{K_{0}},\end{array}$
for which we have $d_{J}\circ\Psi=\Psi\circ d^{\prime}_{J}$. This proves the claim.
Moreover, by a standard argument using a homotopy from $\psi$ to $\operatorname{id}_{T^{*}\mathbb{R}^{3}}$ in $\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$, we can show the invariance of the Floer cohomology under Hamiltonian isotopies, which we state as follows.
Theorem A.1.
If we choose a pair $(H_{0},J_{0})$ as above to define $(CF^{*}(\mathbb{R}^{3},L_{K_{0}};H_{0}),d_{J_{0}})$ (i.e. the case $\psi=\operatorname{id}_{T^{*}\mathbb{R}^{3}}$), then there exists an isomorphism of $\mathbb{Z}/2$-graded $\mathbb{F}_{2}[\lambda^{\pm}]$-modules
$HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H,J))\cong HF^{*}(\mathbb{R}^{3},L_{K_{0}};(H_{0},J_{0})).$
Now, we prove Proposition 2.7 about a clean Lagrangian intersection along a circle.
Proof of Proposition 2.7.
For $\varphi\in\operatorname{Ham}_{c}(T^{*}\mathbb{R}^{3})$ in the statement of Proposition 2.7, let us set $\psi\coloneqq\varphi^{-1}$. Then, we can change the setup as follows: $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ have a clean intersection along $\psi(K_{1})$, where $K_{1}$ is an oriented knot in $\mathbb{R}^{3}$. We may assume that for sufficiently small $\varepsilon>0$, $\psi(L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3})\subset L_{K_{0}}$. We fix an orientation of $\psi(K_{1})$ via $\psi\colon K_{1}\to\psi(K_{1})$ and define $a\in\mathbb{Z}$ by $[\psi(K_{1})]=a[K_{0}]\in H_{1}(L_{K_{0}})$. Let $(-1)^{\sigma}$ denote the sign of orientation of the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$. Then, the claim of the proposition is equivalent to that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$.
We choose a Mose function $f\colon\psi(K_{1})\to\mathbb{R}$ which has only two critical points $x_{0},x_{1}\in\psi(K_{1})$ such that $f$ takes its minimum at $x_{0}$ and its maximum at $x_{1}$. We also choose a Riemannian metric $g$ on $\psi(K_{1})$. By using $g$, we obtain the gradient vector field $V_{f}$ for $f$ by $g(V_{f},\cdot)=df$. Then, we consider the set $\mathcal{T}_{(f,g)}(x_{0},x_{1})$ of trajectories $\gamma\colon\mathbb{R}\to\psi(K_{1})$ such that:
  • $\frac{d\gamma}{ds}(s)=(V_{f})_{\gamma(s)}$ for every $s\in\mathbb{R}$.
  • $\lim_{s\to-\infty}\gamma(s)=x_{0}$ and $\lim_{s\to\infty}\gamma(s)=x_{1}$.
Since $\psi(\mathbb{R}^{3})$ and $L_{K_{0}}$ intersects cleanly along $\psi(K_{1})$ and $\psi(K_{1})$ is connected, we can apply [18, Theorem 3.4.11, Lemma 3.4.12] to get the following results: From the pair $(f,g)$ with $||f||_{C^{1}}$ sufficiently small, we can construct a Hamiltonian $H_{f}\colon T^{*}\mathbb{R}^{3}\to\mathbb{R}$ supported in a neighborhood of $\psi(K_{1})$ and an almost complex structure $J_{g}$ on $T^{*}\mathbb{R}^{3}$ such that the pair $(H_{f},J_{g})$ is regular and the following hold:
  • $P_{H_{f}}\cap L_{K_{0}}=\{x_{0},x_{1}\}$.
  • There is a bijection
    $\Phi\colon\coprod_{A\in H_{1}(L_{K_{0}})}\mathcal{M}_{(H_{f},J_{g}),A}(x_{0},x_{1})\to\mathcal{T}_{(f,g)}(x_{0},x_{1})\colon u\mapsto u(\cdot,1).$
It is clear that $\mathcal{T}_{(f,g)}(x_{0},x_{1})$ has only two elements. Let us label them as $\gamma_{1}$ and $\gamma_{2}$ so that $\frac{d\gamma_{2}}{ds}(s)$ is a positive vector with respect to the orientation of $\psi(K_{1})$. We tale $k\in\mathbb{Z}$ such that $\Phi^{-1}(\gamma_{1})$ belongs to $\mathcal{M}_{(H_{f},J_{g}),k[K_{0}]}(x_{0},x_{1})$. ($k$ depends on the choice of the paths $c_{x_{0}}$ and $c_{x_{1}}$.) Then, $\Phi^{-1}(\gamma_{2})$ belongs to $\mathcal{M}_{(H_{f},J_{g}),A}(x_{0},x_{1})$ for the homology class
$A=k[K_{0}]+[\psi(K_{1})]=(k+a)[K_{0}]$
since $[\psi(K_{1})]\in H_{1}(L_{K_{0}})$ is represented by a loop obtained by concatenating the two paths: $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{1}(-s)$ and $[-\infty,\infty]\to L_{K_{0}}\colon s\mapsto\gamma_{2}(s)$. From this observation, the differential $d_{J_{g}}\colon\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{1}\to\mathbb{F}_{2}[\lambda^{\pm}]\cdot x_{0}$ is given by
$d_{J_{g}}(x_{1})=(\lambda^{k}+\lambda^{k+a})x_{0}.$
Therefore, $HF^{*}(\psi(\mathbb{R}^{3}),L_{K_{0}};(H_{f},J_{g}))$ is computed as follow (we omit writing $(H_{f},J_{g})$):
$\begin{array}[]{lll}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]/(1+\lambda^{a})&\text{ if }a\neq 0,\\ HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}],&HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})\cong\mathbb{F}_{2}[\lambda^{\pm}]&\text{ if }a=0.\end{array}$
Here, $\sigma_{\psi}\in\mathbb{Z}/2$ is the shift of degrees. Note that $\sigma_{\psi}$ depends only on the orientation of $L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}$ induced by the open embedding $\psi\colon L_{K_{1}}\cap D^{*}_{\varepsilon}\mathbb{R}^{3}\to L_{K_{0}}$ and the fixed orientation of $L_{K_{0}}$. As $\mathbb{F}_{2}$-vector spaces,
$\begin{array}[]{lll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=|a|&\text{ if }a\neq 0,\\ \dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\psi}}(\psi(\mathbb{R}^{3}),L_{K_{0}})=\infty&\text{ if }a=0.\end{array}$ (15)
If we consider a special case where $\psi$ is the identity map $\operatorname{id}=\operatorname{id}_{T^{*}\mathbb{R}^{3}}$, then $K_{1}=\psi(K_{1})=K_{0}$ and we have $a=\pm 1$. Therefore,
$\begin{array}[]{ll}\dim_{\mathbb{F}_{2}}HF^{0+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=0,&\dim_{\mathbb{F}_{2}}HF^{1+\sigma_{\operatorname{id}}}(\mathbb{R}^{3},L_{K_{0}})=1.\end{array}$ (17)
From the definition of the orientation sign $(-1)^{\sigma}$, the shift of degrees $\sigma_{\psi}$ is equal to $\sigma_{\operatorname{id}}+\sigma$ modulo $2$. From (15), (17) and Theorem A.1, it follows that $a\in\{1,-1\}$ and $\sigma\equiv 0\mod 2$. ∎